Determine the set of points at which the function is continuous.
1) F(x,y)=sin(xy)÷(e^x-y^2)
2) F(x,y)=x-y÷(1+x^2+y^2)
3) F(x,y)=arctan(x+√y)
4) F(x,y)=e^x^2y+√x+y^2
5) G(x,y)=ln(x^2+y^2-4)
1
Expert's answer
2020-11-25T16:20:24-0500

Given first function

(i) "f(x,y)=\dfrac{sin(xy)}{e^x-y^2}"


"f(x,y)" is continous for all values except at "e^x-y^2=0"

"\implies e^x=y^2"


Taking log on both sides

"\implies xlog e=2logy \\\implies x=2logy"


(ii) "f(x,y)=\dfrac{x-y}{1+x^2+y^2}"

"since" "1+x^2+y^2>0"

"f(x,y)" is continous of alll value of "x" and "y"


(iii) "f(x,y)=arctan(x+\sqrt{y})"

"f(x,y)" is continuous for all value of "x" and "y" except "y<0"


(iv) "f(x,y)=e^{x^2y}+\sqrt{x}+y^2"


Given function is continuous for all values of "x" and "y" except at "x<0"


(v) "f(x,y)=ln(x^2+y^2-4)"


Given function is continous for "x^2+y^2>4"

"\implies x\ge 0, y\ge 2"

So the set of values for which given function are cintinuous

are "x\ge 0,y\ge 0"